OpenStudy (anonymous):

Ginny's class ordered 4 pizzas. They ate seven-twelfths of each pizza. How many pizzas did they eat? tell me

OpenStudy (anonymous):

They ordered 4 pizzas.. they ate seven twelfths so, They ate: \[\frac{7}{12} \times 4 = \frac{7}{3}\]